OpenStudy (anonymous):

can anyone help me solve 5x+10=58=11x+2? please

OpenStudy (ranga):

There is an equal sign on either side of 58?

OpenStudy (anonymous):

no cause its solving eqautions using the exterior angle theorem

OpenStudy (ranga):

Can you post the original problem exactly the way it appears?

OpenStudy (anonymous):

well ive tried but all i can do is tell you where everything goes angle T is 5x+10 angle S is 58 degrees and then Angle U is 11x+2 which is the extended side ...and all it is,is an upside down triangle

OpenStudy (ranga):

So angle T and angle S are inside the triangle and angle U is outside the triangle?

OpenStudy (anonymous):

yupp

OpenStudy (ranga):

exterior angle of a triangle = sum of the opposite interior angles. 11x + 2 = 5x + 10 + 58 11x + 2 = 5x + 68 subtract 2 11x = 5x + 66 subtract 5x 11x - 5x = 66 6x = 66 divide by 6 x = 11

OpenStudy (anonymous):

thanks :)

OpenStudy (ranga):

yw.

OpenStudy (anonymous):

could you help me with 15x+5+22xx+4=120

OpenStudy (ranga):

What is 22xx? Are you typing the problem correctly. In the previous question you had typed two equal signs on either side of 58. The = on the left should have been a + sign and that could have avoided a lot of confusion.

OpenStudy (anonymous):

sorry i meant 18x+5=-1+8x+46

OpenStudy (anonymous):

im sorry i didnt mean to confuse you sorry

OpenStudy (anonymous):

typos

OpenStudy (ranga):

18x+5=-1+8x+46 first simplify the constants on the right: 18x + 5 = 8x + 47 subtract 8x from both sides. Then subtract 5 from both sides solve for x.

OpenStudy (anonymous):

so it should like 18x+5=8x+47

OpenStudy (anonymous):

Simplifying 15x + 5 + 22x + 4 = 120 Reorder the terms: 5 + 4 + 15x + 22x = 120 Combine like terms: 5 + 4 = 9 9 + 15x + 22x = 120 Combine like terms: 15x + 22x = 37x 9 + 37x = 120 Solving 9 + 37x = 120 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-9' to each side of the equation. 9 + -9 + 37x = 120 + -9 Combine like terms: 9 + -9 = 0 0 + 37x = 120 + -9 37x = 120 + -9 Combine like terms: 120 + -9 = 111 37x = 111 Divide each side by '37'. x = 3 Simplifying x=3

OpenStudy (anonymous):

ohh wow lol thats a lot i thought it x was 11

OpenStudy (anonymous):

well the site will help you dont have to ask anyone anymore on these

OpenStudy (anonymous):

hope i helped

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

could you send me the link again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.geteasysolution.com